When a 275 mL sample of nitrogen is kept at constant temperature, it has a pressure of 2.0 atmospheres. What pressure does it exert when compressed to 125 mL?
1
Expert's answer
2020-10-22T08:38:42-0400

According to Boyle's law, at constant temperature pressure is inversely proportional to volume:

"p\propto\frac{1}{V}" .

Using this law, one can calculate the pressure of the system which volume was changed at constant temperature:

"p_2 = \frac{p_1V_1}{V_2} = \frac{2.0\text{ atm}\cdot275\text{ mL}}{125\text{ mL}} = 4.4" atm.

Answer: the sample of nitrogen exerts pressure of 4.4 atm when compressed to 125 mL.